# A Principled Framework for Evaluating on Typologically Diverse Languages

### v2

@@ 1 Introduction
A shared factor behind the success of these models is their reliance on large volumes of textual data. These data-driven approaches are claimed to be language-independent, as they are, in principle, applicable to any language given enough training data. However, language-independent systems are not language agnostic (Bender, 2009, 2011). Current algorithms are designed with an English-centric perspective in mind, while characteristics (such as low morphological complexity) of the English language cannot be assumed to transfer to other languages. To be best of our knowledge, there is currently no systematic investigation of how the properties of languages included in the evaluation influence the performance estimations of multilingual language models. It remains unclear how multilingual these models truly are and to what extent the language-independent assumption really holds against the world’s language diversity.
To assess whether a language model performs well across languages, ideally, one would evaluate it on all languages in the world. However, collecting high-quality data on such a scale is not feasible. Therefore, multilingual models are evaluated on a sample of the world’s languages. To ensure generalizability, such a language sample should be diverse, with varying characteristics and properties. However, existing approaches to this sampling process have not been optimal. Previous work has established that there is no clear terminology or methodological consensus for what constitutes ‘typological diversity’ (Ploeger et al., 2024). Currently, many approaches in NLP use phylogenetic heuristics to ensure ‘typological diversity’. In this paper, we show that this approximation of typological diversity through language phylogeny has severe shortcomings, and we provide a framework for systematically selecting languages based on typological distance measures. Our framework implements two sampling methods that are widely established within the field of linguistic typology. We demonstrate how our framework can be used for diverse typological language sampling, how it can help guide dataset expansions, and how it has use beyond typology.
@@
(i) We establish that language phylogeny is limited when it comes to assessing typological diversity; (ii) we provide a method for quantifying the typological distance between pairs of languages; (iii) we provide a systematic framework for selecting typologically diverse languages for multilingual evaluation scenarios; (iv) we introduce measures of typological diversity, and compare these with existing ones; (v) we show that typological diversity matters in downstream evaluation; (vi) we provide a Python package that facilitates typologically diverse language selection, which is publicly available in the following repository:
@@ 2 Background
Linguistic typology can be described as the study of structural similarities and differences across the world’s languages (Kashyap, 2019). Within linguistic typology, an important research direction is the investigation of general patterns for these similarities and differences. For example, Greenberg (1963) finds that in languages with verb-object word order, adpositions tend to be placed before their objects. For drawing such general conclusions about human languages as a whole, testing linguistic hypotheses on only a handful of related languages is insufficient (Rijkhoff et al., 1993; Guzmán Naranjo and Becker, 2022). Instead, generalizable conclusions in typology require adequate sampling strategies: findings should be supported by evidence across a diverse range of languages. Thus, representative language sampling has long been a central methodological issue in the field of linguistic typology. Given the recent advances in language technology beyond English, language sampling has become increasingly relevant for multilingual NLP. Drawing generalizable conclusions about multilingual model performance requires tests across diverse languages. Moreover, gaining insight into the skews of evaluation language sets may help to identify weaknesses of current applications. In this section, we discuss relevant sampling strategies and terminology from the field of linguistic typology and how they relate to language sampling in NLP.
@@
Bell (1978) introduced central notions for language sampling in typology. Firstly, the sampling universe is the class of objects under study. In typological studies, this could for instance be the set of all possible human languages. In NLP, the sampling universe often corresponds to all existing natural languages that a certain language technology is claimed to generalize to. The sampling frame provides access to the sampling universe. It is the concrete set of languages one can actually sample from. Sjöberg (2023) further describes the distinction between a catalogue frame, which is the set of languages we know to exist, and a corpus frame, which is the set of languages for we have relevant research materials. For our purposes, we will treat the catalogue frame as all languages we have typological information for. We will treat the corpus frame as the set of languages that we have available datasets for. Given the data requirements of most techniques in NLP, the corpus frame is of vital importance. Lastly, the sample is the set of languages that one draws from the frame, with the aim to reflect the properties of the sampling universe. In NLP, this can correspond to the concrete set of languages that one tests or evaluates an application on.
@@
For extracting a sample from the frame, there are multiple popular types of methods in linguistic typology. Rijkhoff and Bakker (1998) divide sampling methods into three categories: random, probability and variety sampling. Each of these sampling methods are relevant for different research questions. Random sampling entails selecting languages without any criteria. Without ‘stratification’, the grouping of the frame before sampling, it is possible that resulting samples are in large part made up of similar languages. As such, these samples could be skewed towards languages from certain phylogenetic or geographical groups. This is especially likely if the sampling frame contains a skew. Such samples, given they are large enough, are commonly used to look into the occurrence frequency of some linguistic phenomenon, but not necessarily for generalizable conclusions about language. In probability sampling, one aims for a sample that contains independent languages. Ideally, the resulting sample should be free of bias. For instance, it should not contain a skew towards one or a few language families. Variety sampling aims at sampling languages such that the linguistic diversity of the world’s languages is captured as much as possible. This is because ‘exceptional types test the rule’ (Perkins, 1988). In the context of NLP, Ponti et al. (2020) write that choosing a variety sample tests the robustness of a language model to unseen typological features, as it includes outliers. These three sampling methods each come with different implications in terms of sample size. Because of the lack of stratification, a random sample ‘must be relatively large in size to be able to produce reliable results’ (Rijkhoff and Bakker, 1998). For variety sampling, it is also often the case that the more languages are included, the better. This is because including more languages means that outliers and uncommon properties are more likely to be captured (Miestamo, Bakker, and Arppe, 2016). However, for probability sampling, there is a trade-off between independence and coverage. The more languages one includes, the more difficult it becomes to preserve independence between these languages. To illustrate: if the number of languages one samples is larger than the number of language families in the frame, then multiple languages from the same family have to be sampled.
@@
Bender (2009, 2011) and Pikuliak and Simko (2022) extensively discussed the need for diverse evaluation language selection, and the potential for linguistic typology to facilitate this. However, to date, there are only a handful of works in NLP that aimed to apply these suggestions. In fact, Ploeger et al. (2024) ascertained that ‘typologically diverse’ sampling methods in NLP are mostly flawed. Firstly, if any stratification criterion is mentioned in NLP, this is usually based on phylogeny (e.g., Yadavalli, Yadavalli, and Tobin, 2023; Ács, Kádár, and Kornai, 2021; Xu et al., 2020). To the best of our knowledge, no approach in NLP uses a tree structure to apply genealogical stratification. Instead, sampling from different families is common. To illustrate: Majewska et al. (2020) ‘sampled languages from 5 different language families to ensure typological diversity’. However, there is no evidence that phylogenetic relations directly imply typological relations (Dahl, 2008). We further explore this in Section 3. Secondly, most sampling stratification is applied post-hoc. Commonly, a set of languages is selected, and only then is it described in terms of diverse language families and sometimes typological diversity. For example, the typology index by Ponti et al. (2020) (see Section 6.2) is only applied after language selection. As such, it does not promote principled language selection. Our work differs from these approaches, in that we perform informed selection, instead of informed analysis.
@@ 3 Why is Phylogeny Insufficient?
All in all, we conclude that sampling with phylogenetic stratification is not ideal for NLP purposes. In linguistic typology, sampling using a proxy is often necessary. Directly using typological values instead is to be avoided, because then using those same variables for sampling introduces circularity. However, in NLP, the variable under investigation is commonly something different, such as multilingual model performance. Thus, there is no circularity with sampling directly from typological features. This avoids many of the issues with strict group sampling: we can take into account granularity, include languages which are typically difficult to place in one of those strict groups, we can control the group size and are not affected by randomness in sampling. Additionally, by sampling through typological features directly, we gain immediate insight into the typological diversity of the sample. In the next Section 5, we present a method that does exactly this.
@@ 5.1 Typological Feature Vectors
Our framework is compatible with any kind of information about languages. The framework can also be used starting from step 3 when there are already pairwise distances available. These language distances are not required to be based on typology.111We show an application of this in Section 8.2. In this work, we specifically use Grambank, for multiple reasons. Firstly, Grambank was developed with computational applications in mind, taking specific care to avoid logical dependencies between feature values (Haynie et al., 2023). Logical dependencies are introduced when the value of one feature implies another. These are problematic for calculating (language) distances, as features that directly imply each other are then weighted more than features that are not logically implied by others.222It should be noted that a lack of logical dependencies does not exclude statistical dependencies, as some feature values may co-occur. Additionally, since our work focuses on text processing specifically, the morphosyntactic domain of Grambank is particularly relevant. By contrast, databases such as WALS also contain phonological features, while phonology is generally separate from text. For our main experiments, we therefore treat Grambank as our catalogue frame. Still, if one aims to conduct a phonological study, one could use phonological features with our framework instead.
@@ 5.3.1 MaxSum Diversity
For our typologically informed variety sampling, we treat the problem as a maximum diversity problem (MDP).444Martí et al. (2013) list a number of alternative names used for the problem: maximum dispersion, max-avg dispersion, -dispersion, -dispersion-sum, edge-weighted clique, remote clique, maximum edge-weighted subgraph, dense -subgraph, -defense, -defence-sum and equitable dispersion. This entails finding a size set of points where the sum of distances between all points in the set is maximal (MaxSum). Approaching the problem as such optimizes for sampling for large total distance, capturing outliers (see Figure 4), which is the objective of variety sampling.
Kuo, Glover, and Dhir (1993) showed that the ‘clique problem’, finding subsets of vertices in a graph that are all adjacent, can be reduced to MDP. Since the clique problem is NP-hard, then so must MDP be. Famously, there is no known efficient algorithm for finding optimal solutions to NP-hard problems. We also experienced that a brute force algorithm did not terminate when run on the dataset. Martí et al. (2013) give an overview of heuristics for MDP and conclude that even simple heuristics give good solutions. We implement one such simple heuristic. Given pairwise language distances, we first find the language that is most distant from all others. For this, we take the language where the sum of distances with all other languages is largest. The motivation behind this is that selecting the language with the most ‘unusual’ typological properties (or combinations thereof) already ensures some outlier. Next, we add the language that is furthest from the first language to the sample. We then sum the distances from these two languages to every other language. The language that has the largest summed distance is added to the selection. We repeat this process until the desired size of the sample is reached. The full procedure is described in Algorithm 1.
@@ 5.3.2 MaxMin Diversity
In sampling with the MaxSum objective, the total typological distance between languages is maximized. Individual outliers within the sample may be close (see Figure 4), which may introduce a skew in a sample.555 The distribution of language types can be visualized as distributed normally around the center, see Dryer (1998). However, for probability sampling, one would want to preserve the independence between languages. To this end, we approach the diversity problem as a MaxMin Diversity Problem, which is a second popular diversity sampling objective (Parreño, Álvarez-Valdés, and Martí, 2021). Instead of optimizing for the maximum total distance, we optimize for maximizing the smallest distance between any two points in the set. Thus, we select languages such that the closest two are maximally typologically distant. This problem is NP hard as well, which means we again need to rely on heuristics. Similar to MaxSum diversity, we implement a simple heuristic. As in the MDP approach, we first select the language that is most distant from all others. Then, we again take the language furthest from that language. Then until the desired sample size () is reached, we add the next language to the sample that retrieves the highest minimum distance to the already selected points. This last step is the key difference between the objectives of both algorithms. The full process is described in Algorithm 2.
@@ 6.1 Baselines
The most popular stratification in NLP, if any, is sampling based on language families. This method can be motivated from the perspective of linguistic typology, where phylogenetic stratification is popular. Contrary to linguistic typology, advanced sampling methods from language trees are not relevant to NLP (Section 2). We implement phylogenetic stratification by sampling uniformly from groupings on two levels: language families666As specified in Glottolog v4.4: 215 families. and genera777As specified in WALS v2020.3: 612 genera.. If the requested sample size is bigger than the number of groupings, we uniformly sample again from the groupings until we reach the requested size. Note that this already more principled than most sampling approaches in NLP. The convenience baseline is indirectly also often influenced or motivated by phylogeny, albeit in less principled manner. As such, our phylogenentic baseline is a ‘best case scenario’ for principled, typologically diverse language sampling using phylogenetic data.
@@ 6.2 Metrics
Similar to the diversity index reported in Ponti et al. (2020), we report the entropy of the feature values that occur in a language sample. This gives us insight into the spread within features. For example, if a sample contains for a given feature , FVI is maximal, but the spread is low. The entropy for is lower than a more diverse sample with values . We take the average of this metric over all features. A difference with previous work is that Ponti et al. (2020) based their entropy calculations based on 103 unnamed typological features from URIEL (Littell et al., 2017). This is not reproducible, as they do not report which typological features were used. Furthermore, as URIEL contains logical dependencies, the included typological features are not necessarily weighted equally. The entropy of a set of languages is the average entropy over all features:

### v3

@@ 1 Introduction
A shared factor behind the success of these models is their reliance on large volumes of textual data. These data-driven approaches are claimed to be language-agnostic, as they are, in principle, applicable to any language given enough training data. However, language-agnostic systems are not language-independent (Bender, 2009, 2011). Current algorithms are designed with an English-centric perspective in mind, while characteristics (such as low morphological complexity) of the English language cannot be assumed to transfer to other languages. To the best of our knowledge, there is currently no systematic investigation of how the properties of languages included in the evaluation influence the performance estimations of multilingual language models. It remains unclear how well these models perform on diverse languages systematically sampled.
To assess whether a language model performs well across languages, ideally, one would evaluate it on all languages in the world. However, collecting high-quality data on such a scale is not feasible. Therefore, multilingual models are evaluated on a sample of the world’s languages. To increase generalizability, such a language sample should be diverse, with varying characteristics and properties. However, existing approaches to this sampling process have not been optimal. Previous work has established that there is no clear terminology or methodological consensus for what constitutes ‘typological diversity’ (Ploeger et al., 2024). Currently, many approaches in NLP use phylogenetic heuristics to ensure ‘typological diversity’. In this paper, we show that this approximation of typological diversity through language phylogeny has severe shortcomings, and we provide a framework for systematically selecting languages based on typological distance measures. Our framework implements two sampling methods that are widely established within the field of linguistic typology. We demonstrate how our framework can be used for diverse typological language sampling, how it can help guide dataset expansions, and how it has use beyond typology.
@@
(i) We establish that language phylogeny is limited when it comes to assessing typological diversity; (ii) we provide a method for quantifying the typological distance between pairs of languages; (iii) we provide a systematic framework for selecting typologically diverse languages for multilingual evaluation scenarios; (iv) we introduce four measures of typological diversity; (v) we show that typological diversity matters in downstream evaluation; (vi) we provide a Python package that facilitates typologically diverse language selection, which is publicly available in the following repository:
@@ 2 Background
Representative language sampling has long been a central methodological issue in the field of linguistic typology. In order to assess its applicability to sampling in NLP, we discuss this line of research in more detail.
Linguistic typology can be described as the study of structural similarities and differences across the world’s languages (Kashyap, 2019). Within linguistic typology, an important research direction is the investigation of general patterns for these similarities and differences. For example, Greenberg (1963) finds that in languages with verb-object word order, adpositions tend to be placed before their objects. For drawing such general conclusions about human languages as a whole, testing linguistic hypotheses on only a handful of related languages is insufficient (Rijkhoff et al., 1993; Guzmán Naranjo and Becker, 2022). Instead, generalizable conclusions in typology require adequate sampling strategies: findings should be supported by evidence across a diverse range of languages. Thus, language sampling has received considerable attention in typological research. Given the recent advances in language technology beyond English, language sampling has become increasingly relevant for multilingual NLP. Drawing generalizable conclusions about the performance of multilingual models requires tests across diverse languages. Moreover, gaining insight into the skews of evaluation language sets may help to identify weaknesses of current applications. In this section, we discuss relevant sampling strategies and terminology from the field of linguistic typology and how they relate to language sampling in NLP, despite differences in objectives..
@@
Bell (1978) introduced central notions for language sampling in typology. To establish a common ground for approaching language sampling more systematically in NLP, we relate these terms to NLP. Firstly, the sampling universe is the class of objects under study. In typological studies, this could for instance be the set of all possible human languages. In NLP, the sampling universe often corresponds to all existing natural languages that a certain language technology is claimed to generalize to. The sampling frame provides access to the sampling universe. It is the concrete set of languages one can actually sample from. Sjöberg (2023) further describes the distinction between a catalogue frame, which is the set of languages we know to exist, and a corpus frame, which is the set of languages for which we have relevant research materials. For our purposes, we will treat the catalogue frame as all languages we have typological information for. We will treat the corpus frame as the set of languages that we have available datasets for. Given the data requirements of most techniques in NLP, the corpus frame is of vital importance. Lastly, the sample is the set of languages that one draws from the frame, with the aim to reflect the properties of the sampling universe. In NLP, this can correspond to the concrete set of languages that one tests or evaluates an application on.
@@
For extracting a sample from the frame, there are multiple popular types of methods in linguistic typology. Rijkhoff and Bakker (1998) divide sampling methods into three categories: random, probability and variety sampling. Each of these sampling methods is relevant for different research questions. Random sampling entails selecting languages without any criteria. Without ‘stratification’, the grouping of the frame before sampling, it is possible that resulting samples are in large part made up of similar languages. As such, these samples could be skewed towards languages from certain phylogenetic or geographical groups. This is especially likely if the sampling frame contains a skew. Such samples, given they are large enough, are commonly used to look into the occurrence frequency of some linguistic phenomenon, but not necessarily for generalizable conclusions about language. In probability sampling, one aims for a sample that contains independent languages. Ideally, the resulting sample should be free of bias. For instance, it should not contain a skew towards one or a few language families. Variety sampling aims at sampling languages such that the linguistic diversity of the world’s languages is captured as much as possible. This is because ‘exceptional types test the rule’ (Perkins, 1988). It should be noted, however, that rather than creating statistical models of all languages of the world, sampling in NLP is rather concerned with testing multilingual model robustness across languages. In the context of NLP, Ponti et al. (2020) write that choosing a variety sample tests the robustness of a language model to unseen typological features, as it includes outliers. These three sampling methods each come with different implications in terms of sample size. Because of the lack of stratification, a random sample ‘must be relatively large in size to be able to produce reliable results’ (Rijkhoff and Bakker, 1998). For variety sampling, it is also often the case that the more languages are included, the better. This is because including more languages means that outliers and uncommon properties are more likely to be captured (Miestamo, Bakker, and Arppe, 2016). However, for probability sampling, there is a trade-off between independence and coverage. The more languages one includes, the more difficult it becomes to preserve independence between these languages. To illustrate: if the number of languages one samples is larger than the number of language families in the frame, then multiple languages from the same family have to be sampled.
@@
Bender (2009, 2011) and Pikuliak and Simko (2022) extensively discussed the need for diverse evaluation language selection, and the potential for linguistic typology to facilitate this. However, to date, there are only a handful of works in NLP that aimed to apply these suggestions. In fact, Ploeger et al. (2024) ascertained that ‘typologically diverse’ sampling methods in NLP are mostly flawed. Firstly, if any stratification criterion is mentioned in NLP, this is usually based on phylogeny (e.g., Yadavalli, Yadavalli, and Tobin, 2023; Ács, Kádár, and Kornai, 2021; Xu et al., 2020). To the best of our knowledge, no approach in NLP uses a tree structure to apply genealogical stratification. Instead, sampling from different families is common. To illustrate: Majewska et al. (2020) ‘sampled languages from 5 different language families to ensure typological diversity’. However, there is no evidence that phylogenetic relations directly imply typological relations (Dahl, 2008). We further explore this in Section 3. Secondly, most sampling stratification is applied post-hoc. Commonly, a set of languages is selected, and only then is it described in terms of diverse language families and sometimes typological diversity. For example, the typology index by Ponti et al. (2020) (see Section 6.2) is only applied after language selection. As such, it does not promote principled language selection. Our work differs from these approaches, in that we perform informed selection, instead of only informed analysis.
@@ 3 Why is Phylogeny Insufficient?
All in all, we conclude that sampling with phylogenetic stratification is not ideal for NLP purposes. In linguistic typology, sampling using a proxy is often necessary. Directly using typological values instead is to be avoided, because then using those same variables for sampling introduces circularity. However, in NLP, the variable under investigation is commonly something different, such as the performance of multilingual models. Thus, there is no circularity with sampling directly from typological features. This avoids many of the issues with strict group sampling: we can take into account granularity, include languages which are typically difficult to place in one of those strict groups, we can control the group size and are not affected by randomness in sampling. Additionally, by sampling through typological features directly, we gain immediate insight into the typological diversity of the sample. In the next Section 5, we present a method that does exactly this.
@@ 5.1 Typological Feature Vectors
Our framework is compatible with any kind of information about languages. The framework can also be used starting from step 3 when there are already pairwise distances available. These language distances are not required to be based on typology.111We show an application of this in Section 8.2. In this work, we specifically use Grambank, for multiple reasons. Firstly, Grambank was developed with computational applications in mind, taking specific care to avoid logical dependencies between feature values (Haynie et al., 2023). Logical dependencies are introduced when the value of one feature implies another. These are problematic for calculating (language) distances, as features that directly imply each other are then weighted more than features that are not logically implied by others.222It should be noted that a lack of logical dependencies does not exclude statistical dependencies, as some feature values may co-occur. Additionally, since our work focuses on text processing specifically, the morphosyntactic domain of Grambank is particularly relevant. For our main experiments, we therefore treat Grambank as our catalogue frame. Still, if one aims to conduct a phonological study, one could use phonological features with our framework instead. It should be noted that the representativeness of this sampling approach heavily depends on the input features. For example, using input features that are only available for a few languages cannot retrieve representative samples more generally. In addition, statistical dependencies between features can bias the sampling results, making it essential to relate the input features to the use case at hand.
@@ 5.3.1 MaxSum Diversity
For our typologically informed variety sampling, we treat the problem as a maximum diversity problem (MDP).444Mart´ı et al. (2013) list a number of alternative names used for the problem: maximum dispersion, max-avg dispersion, -dispersion, -dispersion-sum, edge-weighted clique, remote clique, maximum edge-weighted subgraph, dense -subgraph, -defense, -defence-sum and equitable dispersion. This entails finding a size set of points where the sum of distances between all points in the set is maximal (MaxSum). Approaching the problem as such optimizes for sampling for large total distance, capturing outliers (see Figure 4), which is the objective of variety sampling.
Kuo, Glover, and Dhir (1993) showed that the ‘clique problem’, finding subsets of vertices in a graph that are all adjacent, can be reduced to MDP. Since the clique problem is NP-hard, then so must MDP be. Famously, there is no known efficient algorithm for finding optimal solutions to NP-hard problems. We also experienced that a brute force algorithm did not terminate when run on the dataset. Mart´ı et al. (2013) give an overview of heuristics for MDP and conclude that even simple heuristics give good solutions. We implement one such simple heuristic. Given pairwise language distances, we first find the language that is most distant from all others. For this, we take the language where the sum of distances with all other languages is largest. The motivation behind this is that selecting the language with the most ‘unusual’ typological properties (or combinations thereof) already ensures some outlier. Next, we add the language that is furthest from the first language to the sample. We then sum the distances from these two languages to every other language. The language that has the largest summed distance is added to the selection. We repeat this process until the desired size of the sample is reached. The full procedure is described in Algorithm 1.
@@ 5.3.2 MaxMin Diversity
In sampling with the MaxSum objective, the total typological distance between languages is maximized. Individual outliers within the sample may be close (see Figure 4), which may introduce a skew in a sample.555 The distribution of language types can be visualized as distributed normally around the center, see Dryer (1998). However, for probability sampling, one would want to preserve the independence between languages. To this end, we approach the diversity problem as a MaxMin Diversity Problem, which is a second popular diversity sampling objective (Parreño, Álvarez-Valdés, and Mart´ı, 2021). Instead of optimizing for the maximum total distance, we optimize for maximizing the smallest distance between any two points in the set. Thus, we select languages such that the closest two are maximally typologically distant. This problem is NP hard as well, which means we again need to rely on heuristics. Similar to MaxSum diversity, we implement a simple heuristic. As in the MDP approach, we first select the language that is most distant from all others. Then, we again take the language furthest from that language. Then until the desired sample size () is reached, we add the next language to the sample that retrieves the highest minimum distance to the already selected points. This last step is the key difference between the objectives of both algorithms. The full process is described in Algorithm 2.
@@ 6.1 Baselines
The most popular stratification in NLP, if any, is sampling based on language families. This method can be motivated from the perspective of linguistic typology, where phylogenetic stratification is popular. Contrary to linguistic typology, advanced sampling methods from language trees are not relevant to NLP (Section 2). We implement phylogenetic stratification by sampling uniformly from groupings on two levels: language families666As specified in Glottolog v4.4: 215 families. and genera777As specified in WALS v2020.3: 612 genera.. If the requested sample size is bigger than the number of groupings, we uniformly sample again from the groupings until we reach the requested size. Note that this is already more principled than most sampling approaches in NLP. The convenience baseline is indirectly also often influenced or motivated by phylogeny, albeit in less principled manner. As such, our phylogenetic baseline is a ‘best case scenario’ for principled, typologically diverse language sampling using phylogenetic data.
@@ 6.2 Metrics
Similar to the diversity index reported in Ponti et al. (2020), we report the entropy of the feature values that occur in a language sample. This gives us insight into the spread within features. For example, if a sample contains for a given feature , FVI is maximal, but the spread is low. The entropy for is lower than that of a more diverse sample with values . We take the average of this metric over all features. A difference with previous work is that Ponti et al. (2020) based their entropy calculations on 103 unnamed typological features from URIEL (Littell et al., 2017). This is not reproducible, as they do not report which typological features were used. Furthermore, as URIEL contains logical dependencies, the included typological features are not necessarily weighted equally. The entropy of a set of languages is the average entropy over all features: